OpenStudy (anonymous):

Joel has 2/3 of a large candy bar that he plans to divide into 4 pieces for himself and his three friends. What fraction of the original large candy bar will each person get? A)3/8 B)8/3 C) 1/6 D)1/12

OpenStudy (anonymous):

2/3:4=2/3*1/4=2/12 =1/6